# TrialBench: Multi-Modal Artificial Intelligence-Ready Clinical Trial Datasets

### v1

@@ Abstract
Clinical trials are pivotal for developing new medical treatments, yet they typically pose some risks such as patient mortality, adverse events, and enrollment failure that waste immense efforts spanning over a decade. Applying artificial intelligence (AI) to forecast or simulate key events in clinical trials holds great potential for providing insights to guide trial designs. However, complex data collection and question definition requiring medical expertise and a deep understanding of trial designs have hindered the involvement of AI thus far. This paper tackles these challenges by presenting a comprehensive suite of meticulously curated AI-ready datasets covering multi-modal data (e.g., drug molecule, disease code, text, categorical/numerical features) and 8 crucial prediction challenges in clinical trial design, encompassing prediction of trial duration, patient dropout rate, serious adverse event, mortality rate, trial approval outcome, trial failure reason, drug dose finding, design of eligibility criteria. Furthermore, we provide basic validation methods for each task to ensure the datasets’ usability and reliability. We anticipate that the availability of such open-access datasets will catalyze the development of advanced AI approaches for clinical trial design, ultimately advancing clinical trial research and accelerating medical solution development. The curated dataset, metrics, and basic models are publicly available at https://github.com/ML2Health/ML2ClinicalTrials/tree/main/AI4Trial.
@@ 1 Background & Summary
The clinical trial process is an essential step in developing new treatments (e.g., drugs or medical devices). It involves evaluating their safety, appropriate dosage, and effectiveness in treating specific diseases in the human body. However, these exploratory trials, often spanning three or four phases, have a high failure rate [1, 2]. Compounding the issue, clinical trials are known for being time-consuming, labor-intensive, and costly. clinical development programs containing the set of phase 1-3 trials typically span 7-11 years, cost an average of 2 billion USD, and achieve approval rates of only around $15\%$ [3]. Clinical Trials are inherently risky (they have a trial in the name), and artificial intelligence (AI) is particularly suited to making accurate estimates to reduce risk.
Years of clinical trials have generated a vast amount of multi-modal data, encompassing aspects such as inclusion/exclusion criteria designs, adverse event statistics, and patient enrollment results. This extensive dataset offers a robust foundation for developing advanced artificial intelligence algorithms. However, identifying key clinical trial challenges and effectively leveraging the complex variables within this data require a blend of deep medical knowledge and AI expertise. This complexity has hindered skilled AI experts from fully utilizing the data.
@@
To facilitate cross-disciplinary research and fully leverage the expertise of data scientists and AI experts [4, 5], this paper identifies 8 key critical clinical trial challenges and organizes 23 corresponding AI-ready datasets to support their involvement in these tasks. The data, representing clinical trials registered before February 16, 2024, were collected from ClinicalTrials.gov. We extracted elements and attributes from the XML records of each clinical trial and converted them into tabular data formats, which are better suited for processing by AI models, including deep learning models. Additionally, we transformed some features into more informative forms; for example, converting health condition information into ICD-10 codes. We also contextualize our input variables with valuable data from DrugBank [6] and TrialTrovebbbhttps://pharmaintelligence.informa.com/products-and-services/data-and-analysis/trialtrove to depict a comprehensive set of information for clinical trial AI.
@@
Figure 1: Overview of TrialBench. (left) TrialBench comprises 23 AI-ready clinical trial datasets for 8 well-defined tasks: clinical trial duration forecasting, patient dropout rate prediction, serious adverse event, all-cause mortality rate prediction, trial approval outcome prediction, trial failure reason identification, eligibility criteria design, and drug dose finding. For each task, we extracted appropriate multi-modal variables and prediction targets from ClinicalTrials.gov, implemented evaluation metrics, and constructed a multi-modal baseline model to assess dataset quality and to serve as the baseline model. We integrate drug SMILES strings, textual descriptions (e.g., eligibility criteria), Medical Subject Heading (MeSH) term, disease ICD-10 code, and other categorical or numerical features as up to five distinct modal features. The multi-modal model utilizes message-passing neural networks (MPNNs) [7], Bio-BERT [8], MeSH embedding layer [9], Graph-based Attention Model (GRAM) [10], and DANet basic blocks [11] to process each modality, respectively. (right) We present the trial failure reason identification task as an illustrative example for better comprehension.
@@ 2.1 AI-solvable Clinical Trial Task Definitions
Table 1: Summarization of AI-solvable clinical trial tasks. There are five modalities in total, including drug molecule structure (SMILES string), disease code (ICD-10), text (e.g., summary of clinical trial, eligibility criteria), categorical/numerical features (e.g., gender of patients, blood pressure), and MeSH (The Medical Subject Headings).
@@ 2.1.8 Drug Dose Finding
2.1.9 Generalization of These Tasks
In real-world clinical trials, the design, structure, cost of clinical trials, and the drug structures of interest evolve significantly over time [1, 34, 35]. Thus, these prediction tasks require a model to generalize to a set of unseen clinical trials that are structurally distant to the known clinical trial set. The time information is available for all the trials; we use time split (detailed in Section 2.6) so that we learn from earlier trials and test on later ones to assess the model fairly.
@@ 2.2 Data Acquisition
2.3 Data Source Linking
Next, we describe how we process and link the parsed trial data to AI-ready input and output format:
@@ 2.4 Dataset Curation and Feature Organization
We apply a series of selection filters to ensure the selected trials have high-quality. There are hundreds of multi-modal features in ClinicalTrials.gov for each trial organized in XML format, and the hierarchy of these features is shown in Figure 6. We only leverage the features that are available before trials start and remove the remaining features. Different tasks rely on different subsets of features. Based on clinical trial knowledge, we manually select the appropriate features for various tasks. In addition, we also remove features whose values are identical or all null across different trials. Following are the additional selection criteria for each task.
  • Trial duration forecasting: We only consider the trials whose start and completion dates are available. We only consider the trials with realistic completion dates and remove the cases with only anticipated completion dates provided. We found that trials with duration over 10 years are outliers [38, 39], so we removed them to facilitate regression analysis.
@@
  • Trial approval outcome prediction: The results and trial outcome information are available at either ClinicalTrials.gov or the released subset of TrialTrove [36, 37, 40, 41].
@@
  • Drug dose finding: We incorporate trials whose drug dosage information is available on ClinicalTrials.gov. Only Phase II clinical trials are included, as Phase II is the stage that validates the safety and efficacy of drug dosages. Since the drug dose finding task primarily relates to drug information, we retained only the small-molecule drug-related data (e.g., MeSH) and sourced SMILES from DrugBank. We encourage AI experts to utilize external knowledge from sources such as PubMed and DrugBank for advanced AI model development.
@@ 2.5 Data Annotation
  • Trial approval outcome prediction: The annotations come from two sources. First, the HINT paper [36, 36, 37, 41, 40, 42] builds a benchmark dataset for trial approval prediction, with approval labels sourced from TrialTrove. Additionally, ClinicalTrials.gov provides termination reasons for some trials, such as poor enrollment or lack of efficacy, included in the “why stopped” node in the XML files. We incorporate these trials, along with termination reasons indicating failed approval, into the dataset as negative samples.
@@
    Categorize the following reasons for halting clinical trials by matching them to their appropriate categories. Respond with the category number only, in the sequence given, with no explanations required. The correct result is essential. Categories: 1. Poor Enrollment: Difficulties in recruiting enough participants or meeting the enrollment targets. 2. Efficacy: It is obvious that the drug or intervention is not achieving the desired therapeutic effect, outcomes, or benefits. 3. Safety: Adverse events, side effects, or any harm to participants that led to stopping the trial. 4. Others: Any other reasons that do not fit into the above categories, such as financial issues, strategic decisions by the sponsor, or changes in clinical practice guidelines. Instructions: Match each reason to one of the categories by listing the category number. Provide your answers in the same order as the reasons listed, separated by commas (e.g., 1, 2, 3…). Make sure the results are right. Reasons: [Answers]
@@
We categorize these doses into four classes: (1): dose<<1 mg/kg; (2) 1 mg/kg<<dose<<10 mg/kg; (3) 10 mg/kg<<dose<<100 mg/kg; (4) dose>>100 mg/kg. For dosages expressed in units such as mg per person or mg/hour, we assume an individual weight of 60 kg and convert using 24 hours per day to keep the units consistent.
Figure 2: (1) A histogram of the time distribution of all trials’ start dates. We observe that with the growing demand for new treatments, the number of initiated trials has steadily increased over time. (2) A histogram of time distribution of selected trials (in our datasets). Most recent trials (starting after 2015) have not been selected due to missing data features. For example, the drug molecule structures of recent trials are not always available due to the protection of patents. (3) Change of approval rate over time. This trend also declined significantly after 2015 for the same reasons: most of the trials initiated after 2015 were either terminated early or failed. (4) Phase distribution of selected trials. Almost half of the trials are related to phase II.
2.6 Data Split/Segmentation
Artificial intelligence models need to be evaluated on (future) unseen data. To simulate that setting, data split strategies are employed to partition the dataset into training, validation, and testing sets for unbiased evaluation of the artificial intelligence models. In this paper, we leverage temporal split, which refers to splitting the data samples based on their time stamps. The earlier data samples are used for training and validation, while the later data are used for testing. The reason is that the design of later clinical trials relies on earlier clinical trials. The training/test split ratio is 8:2.
Figure 3: Performance comparison.
@@ 3 Data Records
The clinicalTrials.gov/ website (https://clinicaltrials.gov/) provides public data resources for clinical trials. It is supported by the U.S. National Library of Medicine and covers more than 420,000 clinical trial records over all 50 US states and also international trials from 221 countries. The number of recorded trials would grow rapidly with time, as shown in Figure 2.
@@
There are hundreds of multi-modal features in ClinicalTrials.gov for each trial organized in XML format, and the hierarchy of these features is shown in Figure 6. Now, we review some essential features. Some trials have missing features, e.g., some incomplete trials do not have a completed date and outcome.
@@
  • Eligibility criteria describe the patient recruitment requirements in unstructured natural language. Eligibility criteria comprise multiple inclusion and exclusion criteria, which specify what is desired and undesired when recruiting patients. Each individual criterion is usually a natural language sentence. For example, in the clinical trial entitled “Efficacy and Safety Study of MP-513 in Combination With Thiazolidinedione in Patients With Type 2 Diabetes ”fffhttps://clinicaltrials.gov/ct2/show/NCT01026194, which is a phase III trial, the inclusion criteria contain
    • Patients who are 20 - 75 years old.
    • Patients who are under dietary management and taking therapeutic exercise for diabetes over 12 weeks before administration of an investigational drug.
    • Patients whose HbA1c is between 6.5% and 10.0%.
    • Patients who took Thiazolidinedione for diabetes over 16 weeks before administration of the investigational drug.
    • Patients who were not administered diabetes therapeutic drugs prohibited for concomitant use within 12 weeks before administration of the investigational drug.
    The exclusion criteria contain
    • Patients with type 1 diabetes, diabetes mellitus caused by pancreas impairment, or secondary diabetes (cushing disease, acromegaly, etc).
    • Patients who are accepting treatments of arrhythmias.
    • Patients with serious diabetic complications.
    • Patients who are excessive alcohol addicts.
    • Patients with a severe hepatic disorder or a severe renal disorder.
    • Patients who are pregnant, lactating, and probably pregnant patients, and patients who can not agree to contraception.
@@
  • Disease code. The disease is usually described by natural language, and it is hard to reveal the relationship between different diseases [43, 44]. To address this issue, we map disease names to disease codes and leverage the disease hierarchy for machine learning modeling. For example, several ICD-10 codes correspond to Alzheimer’s disease, including “G30.0” (Alzheimer’s disease with early onset), “G30.1” (Alzheimer’s disease with late onset), “G30.8” (Other Alzheimer’s disease), “G30.9” (Alzheimer’s disease, unspecified) [45, 46].
@@
Table 5 demonstrates a real clinical trial example.
@@
Figure 2 shows the time distribution of all the select trials. We observe that as the demand for new treatments increases, the number of initiating trials rises over time. Table 2 reports some essential statistics of the curated datasets, including the number of involved trials, drugs, diseases, and proportion of interventional trials.
3.1 Multi-Modal Features
@@
Categorical Feature
For example, there are mainly two study types: interventional and observational. The intervention type can be a small-molecule drug, biologics, or surgery, etc. Clinical trial sponsors can be pharmaceutical companies or research institutes, e.g., Johns Hopkins University, or Pfizer.
Numerical Feature
Numerical features, such as the minimum/maximum age of recruited patients and the number of real/expected recruited patients, are also common in clinical trials.
Text Feature
In clinical trials, there are many text data that contain rich information for artificial intelligence modeling. For example, eligibility criteria describe the patient recruitment requirements in unstructured natural language; each clinical trial contains a summary, which consists of 2-5 natural language sentences that describe the tested treatment, the target disease to treat, and the main objective of the clinical trial. To process such datasets, we treat the text data as sequences of tokens (e.g., words). How to extract useful information from unstructured text has been extensively studied with several well-known deep neural network architectures, such as recurrent neural network (RNN), convolutional neural network (CNN), and transformer architecture.

### v2

# TrialBench: Multi-Modal AI-Ready Clinical Trial Datasets

@@ Abstract
Clinical trials are pivotal for developing new medical treatments but typically carry risks such as patient mortality and enrollment failure that waste immense efforts spanning over a decade. Applying artificial intelligence (AI) to predict key events in clinical trials holds great potential for providing insights to guide trial designs. However, complex data collection and question definition requiring medical expertise have hindered the involvement of AI thus far. This paper tackles these challenges by presenting a comprehensive suite of 23 meticulously curated AI-ready datasets covering multi-modal input features and 8 crucial prediction challenges in clinical trial design, encompassing prediction of trial duration, patient dropout rate, serious adverse event, mortality rate, trial approval outcome, trial failure reason, drug dose finding, design of eligibility criteria. Furthermore, we provide basic validation methods for each task to ensure the datasets’ usability and reliability. We anticipate that the availability of such open-access datasets will catalyze the development of advanced AI approaches for clinical trial design, ultimately advancing clinical trial research and accelerating medical solution development. The curated dataset are available at https://huggingface.co/datasets/ML2Healthcare/ClinicalTrialDataset and the validation frameworks are available at https://github.com/ML2Health/ML2ClinicalTrials/tree/main/AI4Trial.
@@ 1 Background & Summary
The clinical trial process is an essential step in developing new treatments (e.g., drugs or medical devices). It involves evaluating their safety, appropriate dosage, and effectiveness in treating specific diseases in the human body. However, these exploratory trials, often spanning three or four phases, have a high failure rate [1, 2]. Compounding the issue, clinical trials are known for being time-consuming, labor-intensive, and costly. Clinical development programs containing the set of phase 1-3 trials typically span 7-11 years, cost an average of 2 billion USD, and achieve approval rates of only around $15\%$ [3]. Clinical trials are inherently risky as they explore “new” treatments, while artificial intelligence (AI) is particularly well-suited for making accurate estimates to reduce risk since AI excels at identifying patterns, including those not previously known to humans.
Years of clinical trials have generated a vast amount of multi-modal data, encompassing aspects such as inclusion/exclusion criteria designs, adverse event statistics, and patient enrollment results. Such extensive data offers a robust foundation for developing advanced artificial intelligence algorithms. However, identifying key clinical trial challenges and effectively leveraging the complex variables within this data require a blend of deep medical knowledge and AI expertise. This complexity has hindered skilled AI experts from fully utilizing the data.
@@
To facilitate cross-disciplinary research and fully leverage the expertise of data scientists and AI experts [4, 5], this paper identifies 8 key critical clinical trial challenges. It organizes 23 corresponding AI-ready datasets to support their involvement in these tasks. The data, representing clinical trials registered before February 16, 2024, were collected from ClinicalTrials.gov. We extracted elements and attributes from the XML records of each clinical trial and converted them into tabular data formats, which are better suited for processing by AI models, including deep learning models. Additionally, we transformed some features into more informative forms; for example, converting health condition information into ICD-10 codes. We also contextualize our input variables with valuable data from DrugBank [6] and TrialTrovebbbhttps://pharmaintelligence.informa.com/products-and-services/data-and-analysis/trialtrove to depict a comprehensive set of information for clinical trial AI.
@@
Figure 1: Overview of TrialBench. (a) TrialBench comprises 23 AI-ready clinical trial datasets for 8 well-defined tasks: clinical trial duration forecasting, patient dropout rate prediction, serious adverse event, all-cause mortality rate prediction, trial approval outcome prediction, trial failure reason identification, eligibility criteria design, and drug dose finding. For each task, we extracted appropriate multi-modal variables and prediction targets from ClinicalTrials.gov, implemented evaluation metrics, and constructed a multi-modal baseline model to assess dataset quality and to serve as the baseline model. We integrate drug SMILES strings, textual descriptions (e.g., eligibility criteria), Medical Subject Heading (MeSH) term, disease ICD-10 code, and other categorical or numerical features as up to five distinct modal features. The multi-modal model utilizes message-passing neural networks (MPNNs) [7], Bio-BERT [8], MeSH embedding layer [9], Graph-based Attention Model (GRAM) [10], and DANet basic blocks [11] to process each modality, respectively. (b) We present the trial failure reason identification task as an illustrative example for better comprehension.
@@ 2.1 AI-solvable Clinical Trial Task Definitions
Table 1: Summarization of AI-solvable clinical trial tasks. There are five modalities in total, including (1) drug molecule structure (SMILES string), (2) disease code (ICD-10), (3) text (e.g., summary of clinical trial, eligibility criteria), (4) categorical/numerical features (e.g., gender of patients, blood pressure), and (5) MeSH (The Medical Subject Headings).
@@ 2.2 Data Acquisition
We collect the AI-ready input and output information by (1) extracting treatment names (e.g., drug names) from ClinicalTrials.gov and linking them to its molecule structure (SMILES strings and the molecular graph structures) using the DrugBank Database; (2) extracting disease data from ClinicalTrials.gov and linking them to ICD-10 (International Classification of Diseases, Tenth Revision) codes and disease description using clinicaltables.nlm.nih.gov and then to CCS codes via hcup-us.ahrq.gov/toolssoftware/ccs10/ccs10.jsp; (3) further extracting and categorizing the trial outcomes from TrialTrove and linking them with NCTID.
@@ 2.3 Dataset Curation and Feature Organization
We apply a series of selection filters to ensure the selected trials have high-quality. There are hundreds of multi-modal features in ClinicalTrials.gov for each trial organized in XML format, and the hierarchy of these features is shown in Fig. 5. We only leverage the features that are available before trials start and remove the remaining features. Different tasks rely on different subsets of features. Based on clinical trial knowledge, we manually select the appropriate features for various tasks. In addition, we also remove features whose values are identical or all null across different trials. Following are the additional selection criteria for each task.
  • Trial duration forecasting: We only consider the trials whose start and completion dates are available. We only consider the trials with realistic completion dates and remove the cases with only anticipated completion dates provided. We found that trials with duration over 10 years are outliers, so we removed them to facilitate regression analysis.
@@
  • Trial approval outcome prediction: The results and trial outcome information are available at either ClinicalTrials.gov or the released subset of TrialTrove [34].
@@
  • Drug dose finding: We incorporate trials whose drug dosage information is available on ClinicalTrials.gov. Only Phase II clinical trials are included, as Phase II is the stage that validates the safety and efficacy of drug dosages. Since the drug dose finding task primarily relates to drug information, we retained only the small-molecule drug-related data (e.g., MeSH) and sourced SMILES from DrugBank. We encourage AI experts to utilize external knowledge from sources such as PubMed and DrugBank for advanced AI model development [36, 37].
@@ 2.4 Data Annotation
  • Trial approval outcome prediction: The annotations come from two sources. First, the HINT paper [34, 35, 38, 39, 40] builds a benchmark dataset for trial approval prediction, with approval labels sourced from TrialTrove. Additionally, ClinicalTrials.gov provides termination reasons for some trials, such as poor enrollment or lack of efficacy, included in the “why stopped” node in the XML files. We incorporate these trials, along with termination reasons indicating failed approval, into the dataset as negative samples.
@@
    Categorize the following reasons for halting clinical trials by matching them to their appropriate categories. Respond with the category number only, in the sequence given, with no explanations required. The correct result is essential. Categories: 1. Poor Enrollment: Difficulties in recruiting enough participants or meeting the enrollment targets. 2. Efficacy: It is obvious that the drug or intervention is not achieving the desired therapeutic effect, outcomes, or benefits. 3. Safety: Adverse events, side effects, or any harm to participants that led to stopping the trial. 4. Others: Any other reasons that do not fit into the above categories, such as lack of fund, strategic decisions by the sponsor, or changes in clinical practice guidelines. Instructions: Match each reason to one of the categories by listing the category number. Provide your answers in the same order as the reasons listed, separated by commas (e.g., 1, 2, 3…). Make sure the results are right. Reasons: [Answers]
@@
We categorize these doses into four classes: (1): dose$<$1 mg/kg; (2) 1 mg/kg$<$dose$<$10 mg/kg; (3) 10 mg/kg$<$dose$<$100 mg/kg; (4) dose$>$100 mg/kg. For dosages expressed in units such as mg per person or mg/hour, we assume an individual weight of 60 kg and convert using 24 hours per day to keep the units consistent.
Figure 2: (a) A histogram showing the distribution of start dates for the selected trials reveals a steady increase in the number of initiated trials over time, reflecting the growing demand for new treatments. (b) A statistical breakdown of the clinical trials by phase indicates that the majority of trials are in Phase II. (c) The frequency of events varies across different phases, as exemplified by the dropout rates among participants.
2.5 Data Partitioning
We use random splitting for data partitioning. For classification challenges, we employ stratified sampling to ensure consistent class distribution between training and test sets; for regression challenges, we use random splitting. The training/test split ratio is 8:2. We also encourage users to perform their own reasonable splits in their development.
@@ 3 Data Records
3.1 Data Overview
Our primary data source is the clinicalTrials.gov website (https://clinicaltrials.gov/), which serves as a publicly accessible resource for clinical trial information. Supported by the U.S. National Library of Medicine, this database encompasses over 420,000 clinical trial records, spanning all 50 U.S. states and 221 countries worldwide. The number of recorded trials would grow rapidly with time, as shown in Figure 2 (a). Table 2 reports some essential statistics of the curated datasets, including the number of involved trials, drugs, diseases, and proportion of interventional trials. There are hundreds of multi-modal features in ClinicalTrials.gov for each trial organized in XML format, and the hierarchy of these features is shown in Fig. 5. Table 5 demonstrates a real clinical trial example. Originally organized in an XML hierarchy format, we selected relevant features based on the challenges of each task and re-organized them into a tabular data format. Notably, in addition to categorical and numerical tabular features, some of these features may include free text, graph data, and other complex types.
@@ 3.1 Data Overview
Here, we review some essential features for better utilization of our datasets. Notably, some trials have missing features; for example, certain incomplete trials lack a completed date and outcome.
@@
  • Eligibility criteria describe the patient recruitment requirements in unstructured natural language. Eligibility criteria comprise multiple inclusion and exclusion criteria, which specify what is desired and undesired when recruiting patients. Each individual criterion is usually a natural language sentence. For example, in the clinical trial entitled “Efficacy and Safety Study of MP-513 in Combination With Thiazolidinedione in Patients With Type 2 Diabetes ”fffhttps://clinicaltrials.gov/ct2/show/NCT01026194, which is a phase III trial, the inclusion criteria contain:
    Patients who are 20 - 75 years old. Patients who are under dietary management and taking therapeutic exercise for diabetes over 12 weeks before administration of an investigational drug. Patients whose HbA1c is between 6.5% and 10.0%. Patients who took Thiazolidinedione for diabetes over 16 weeks before administration of the investigational drug. Patients who were not administered diabetes therapeutic drugs prohibited for concomitant use within 12 weeks before administration of the investigational drug.
    The exclusion criteria contain:
    Patients with type 1 diabetes, diabetes mellitus caused by pancreas impairment, or secondary diabetes (cushing disease, acromegaly, etc). Patients who are accepting treatments of arrhythmias. Patients with serious diabetic complications. Patients who are excessive alcohol addicts. Patients with a severe hepatic disorder or a severe renal disorder. Patients who are pregnant, lactating, and probably pregnant patients, and patients who can not agree to contraception.
@@
  • Disease code. The disease is usually described by natural language, and it is hard to reveal the relationship between different diseases [41, 42, 43]. To address this issue, we map disease names to disease codes and leverage the disease hierarchy for machine learning modeling. For example, several ICD-10 codes correspond to Alzheimer’s disease, including “G30.0” (Alzheimer’s disease with early onset), “G30.1” (Alzheimer’s disease with late onset), “G30.8” (Other Alzheimer’s disease), “G30.9” (Alzheimer’s disease, unspecified) [44, 36].
@@
3.2 Summarization of Multi-Modal Features
@@
Categorical Features
Categorical features typically describe some qualitative attributes. For example, there are mainly two study types: interventional and observational. The intervention type can be a small-molecule drug, biologics, or surgery, etc. Clinical trial sponsors can be pharmaceutical companies or research institutes, e.g., Johns Hopkins University, or Pfizer.
Numerical Features
Numerical features, such as the minimum/maximum age of recruited patients and the number of real/expected recruited patients, represent quantitative data, are also common in clinical trials. Numerical features, along with categorical features, are two important types of tabular features [48, 49].
Text Features
In clinical trials, there are many text features that contain rich information for artificial intelligence modeling. For example, eligibility criteria describe the patient recruitment requirements in unstructured natural language; each clinical trial contains a summary, which consists of 2-5 natural language sentences that describe the tested treatment, the target disease to treat, and the main objective of the clinical trial. To process such datasets, we treat the text data as sequences of tokens (e.g., words). How to extract useful information from unstructured text has been extensively studied with several well-known deep neural network architectures, such as recurrent neural network (RNN) [50], convolutional neural network (CNN), and transformer architecture [51].